OpenStudy (anonymous):

Can someone help me with this please. A certain car costs $6,595 before taxes are added. Taxes are $460 and license tags cost $55. What is the overall tax rate (to the nearest tenth)? _____ %

OpenStudy (anonymous):

To calculate any percentage, divide the amount that is the percentage (the tax rate in this case) by the total and then multiply by 100.

OpenStudy (anonymous):

(In this case, the total would be the cost of the car.)

OpenStudy (karatechopper):

Do i need to add the 55 to the carr amount? And then you want me to find tax? or tax amount with just car?

OpenStudy (anonymous):

6,595+460+55 do you add everything first ????

OpenStudy (anonymous):

im just so confused

OpenStudy (anonymous):

It's a confusing question the way it's phrased. When it says that the cost of the car is $6,595 before tax, that's your hint that $6,595 is the amount that will be taxed. The license tags are not taxed (the confusing part). So you have $460 tax on $6,595. If we call the tax rate T (a percentage), the formula for the amount of tax is: \[\frac{T}{100} \times 6595 = 460\] Let me know if you need help from there.

OpenStudy (anonymous):

i think i need help

OpenStudy (anonymous):

@ChristianGeek

OpenStudy (anonymous):

do you multiply 100 in both sides

OpenStudy (anonymous):

np...let's take it to the next step by multiplying both sides by 100 to get rid of the fraction: \[T \times 6595 = 460 \times 100\] Now we can divide both sides by 6595 to isolate T: \[T = \frac{46000}{6595}\] Now you just need a calculator to finish up!

OpenStudy (anonymous):

Yes (to multiplying by 100)...nice job!

OpenStudy (anonymous):

6.9

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

Yes, except since the full fraction is 6.97... you would round up to 7.0

OpenStudy (anonymous):

You can test your answer by plugging it back into the formula.

OpenStudy (anonymous):

they got answer 7.8 how

OpenStudy (anonymous):

Give me a moment...

OpenStudy (anonymous):

if all the numbers you gave are correct then there is no way they could have come up with 7.8!

OpenStudy (anonymous):

(Unless license tags are considered tax.)

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

yes 55 is included

OpenStudy (anonymous):

i just did it thesame way and got 7.8 with 55 included

OpenStudy (anonymous):

Yep, that's what they did. That is extremely confusing. Here's the math: \[\frac{T}{100}\times6595 = 460 + 55\]\[\frac{T}{100}\times6595 = 515\]\[T\times6595 = 515 \times 100\]\[T = \frac{51500}{6595}\] Sorry about that. The problem is definitely not clear.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

My pleasure.